# Describe-and-Dissect: Interpreting Neurons in Vision Networks with Language Models

### v1

@@ Appendix A Appendix
In the appendix, we discuss specifics of the methods and experiments presented in the paper and provide additional results. Appendix Section A.1 provides further details on the method pipeline, Section A.2 presents additional qualitative evaluations, Section A.3 shows results for quantitative comparison to MILANNOTATIONS, Section A.4 provides detailed ablation studies on each step of the DnD pipeline, Section A.5 shows an extension of DnD to return multiple concept labels, and Section A.6 discusses the limitations and directions for future work.
@@ A.1.3 Scoring Function Details
      A.1.4 Selecting a Scoring Function
@@
      A.1.5 Amazon Mechanical Turk Setup
@@ A.4.4 Effects of GPT Concept Summarization
      A.4.5 Effects of Concept Selection
@@ A.5 Multiple Labels
    A.6 Limitations
@@ A.1.1 Detailed Description of Attention Cropping
The attention cropping algorithm described in Sec 3.1 is composed of three primary procedures:
@@
  • Step 1. Global Thresholding using Otsu’s Method. For an activation map of neuron $n$, we define regions with high activations as the ”foreground” and regions with low activations as the ”background”. Otsu’s Method (Otsu, 1979) automatically calculates threshold values which maximizes interclass variance between foreground and background pixels. Interclass variance $\sigma^{2}_{B}$ is defined by $\sigma^{2}_{B}=W_{b}W_{f}(\mu_{b}-\mu_{f})^{2}$, where $W_{b,f}$ denotes weights of background and foreground pixels and $\mu_{b,f}$ denotes the mean intensity of background and foreground pixels respectively. For global thresholding used in DnD, $W_{f}$ is the fraction of pixels in activation map $M_{n}(x_{i})$ above potential threshold $\lambda$ while $W_{b}$ is the fraction of pixels below $\lambda$. A value of $\lambda$ that maximizes $\sigma^{2}_{B}$ is used as the global threshold.
@@
  • Step 3. Crop Activating Images. Bounding boxes recorded from Step 2 are overlaid on corresponding activating images from $D_{probe}$, which are cropped to form $D_{cropped}$. To prevent overlapping attention crops, the IoU score between every crop in $D_{cropped}$ is less than an empirically set parameter $\eta$, where high $\eta$ values yield attention crops with greater similarities.
@@
Figure 4: Detailed Visualization of Attention Cropping Pipeline. All three steps of attention cropping are shown for Layer 2 Neuron 165. Steps 1 and 2 illustrate the derivation of bounding boxes from salient regions in the original activation map and are overlaid on original activating images from $\mathcal{D}_{probe}$ in Step 3. Cropped images are added back to $\mathcal{D}_{probe}$ to form $\mathcal{D}_{probe}\bigcup\mathcal{D}_{cropped}$.
A.1.2 GPT Prompt
The Concept Selection process utilizes GPT to identify and coherently summarize similarities between image captions of every image in $I$. Due to the generative nature of DnD, engineering an precise prompt is crucial to the method’s overall performance. The use of GPT in DnD can be split into two primary cases, each with an individual prompt: 1) Summarizing image captions to better bring out the underlying concepts, and 2) Detecting similarities between the image captions. We explain each part separately.
@@ A.1.2 GPT Prompt
”Only state your answer without a period and quotation marks. Do not number your answer. State one coherent and concise concept label that simplifies the following description and deletes any unnecessary details:”
@@
”Only state your answer without a period and quotation marks and do not simply repeat the descriptions. State one coherent and concise concept label that is 1-5 words long and can semantically summarize and represent most, not necessarily all, of the conceptual similarities in the following descriptions:”
In addition to the prompt, we use few shot-prompting to feed GPT two sets of example descriptions and along with the respective human-identified similarities:
Example 1: Human-Identified Similarity: ”multicolored textiles” Image Captions: ”a purple background with a very soft texture” ”a brown background with a diagonal pattern of lines and lines” ”a white windmill with a red door and a red door in the middle of the picture” ”a beige background with a rough texture of linen” ”a beige background with a rough texture and a very soft texture”
Example 2: Human-Identified Similarity: ”red-themed scenes” Image Captions: ”a little girl is sitting in a red tractor with the word sofy on the front” ”a toy car sits on a red ottoman in a play room” ”a red dress with silver studs and a silver belt” ”a red chevrolet camaro is on display at a car show” ”a red spool of a cable with the word red on it”
A.1.3 Scoring Function Details
Though scoring functions all seek to accomplish the same purpose, the logic behind them can vary greatly and have different characteristics. Simplest scoring functions, such as mean, can be easily skewed by outliers in $\mathcal{H}_{j}$, resulting in final concepts that are loosely related to the features detected by neurons. In this section, we discuss 3 different scoring functions and a combination of them which we experimented with in section A.1.4. The concept with highest score is chosen as the final description for neuron $n$.
@@ A.1.3 Scoring Function Details
[…]
  • Image Products. Let set $\mathcal{D}_{j}^{t}\subset\mathcal{D}_{j}$, such that it keeps the $t$ highest activating images from $\mathcal{D}_{j}$. From the original set of activating images $I$ and $\mathcal{D}_{j}^{t}$, the Image Product score is defined as the average cosine similarity between original highly activating images and the generated images for the concept $j$. We measure this similarity using CLIP-ViT-B/16 (Radford et al., 2021) as our image encoder $E(\cdot)$:
[…]
    See Figure 5 for an illustration of Image Product. Intuitively, Image Products selects the candidate concept whose generated images are most similar to the original highly activating images. However, Image Product doesn’t really account for how highly the new images actually activate the target neuron, which is why we chose to use this method to supplement other scoring functions. The enhanced scoring method, TopK Squared + Image Products, is described below.
[…]
@@
A.1.4 Selecting a Scoring Function
We again used ResNet-50 with Imagenet + Broden as the probing dataset. 50 neurons were randomly chosen from each of the 4 hidden layers, with each neuron evaluated twice, again rating the quality of descriptions on a scale 1-5. The participants in this experiment were volunteers with no knowledge of which descriptions were provided by which methods. Table 6 shows the performance of different scoring functions described in section A.1.3. We observe that all three scoring functions presented perform similarly well, and shows our algorithm is robust and not dependent on the choice of scoring function.
@@ A.1.4 Selecting a Scoring Function
A.1.5 Amazon Mechanical Turk Setup
In this section we explain the specifics of our Amazon Mechanical Turk experiment setup in section 4.3. In the interface shown in figure 6, we display the top 10 highly activating images for a neuron of interest and prompt workers to answer the question: ”Does description accurately describe most of the above 10 images?”. Each given description is rated on a 1-5 scale with 1 representing ”Strongly Disagree”, 3 representing ”Neither Agree nor Disagree”, and 5 representing ”Strongly Agree”. Evaluators are finally asked to select the label that best describes the set of 10 images.
We used the final question to ensure the validity of participants’ assessments by ensuring their choice for best description is consistent with description ratings, If the label selected as best description in the final question was not (one of) the highest rated descriptions by the user, we deemed a response invalid. In our analysis we only kept valid responses, which amounted to 1744/2400 total responses for Resnet-50 and 443/600 responses for ResNet-18, so around 75% of the responses were valid.
@@ A.1.5 Amazon Mechanical Turk Setup
One downside of our experimental setup is that it is not clear which highly activating images we should display to the users and how to display them, and these choices may be important to the final results. In our experiments, we displayed the images in $\mathcal{D}_{\text{probe}}$ with the highest mean activation as done by CLIP-Dissect(Oikarinen & Weng, 2023), but displaying images with highest max activation would likely be better for MILAN (Hernandez et al., 2022). Unfortunately there is no simple answer to which images we should display that would be good for all methods. See Limitations A.6 for more discussion.
@@ A.2 Qualitative Evaluation
Figures 7, 8, 9, and 10 provide supplementary qualitative results for DnD along with results produced by baseline methods for ResNet-50 and ResNet-18. Figure 11 showcases our descriptions on random neurons of the ViT-B/16 Vision Transformer trained on ImageNet, showing good descriptions overall, and higher quality than CLIP-Dissect descriptions. Finally in Figures 12 and 13 we dissect ResNet-50 using the CIFAR-100 training image set as the probing set, showcasing our method still perfoms well with a different probing dataset.
@@ A.3 Quantitative Results: MILANNOTATIONS
In this section we discuss using the MILANNOTATIONS dataset (Hernandez et al., 2022) as a set of ground truths to calculate quantitative results on the intermediate layers of ResNet-152. Table 7 displays an experiment done on 103 randomly chosen ”reliable” neurons across the 4 intermediate layers of ResNet-152. A neuron was deemed ”reliable” if, out of its three corresponding MILANNOTATIONS, two had a CLIP cosine similarity exceeding a certain threshold (for the purposes of this experiment we set that threshold to 0.81). Additionally, we define our summary function $g$ to be spatial max, as this is what was used to calculate the highest activating images for the MILANNOTATIONS. We compare against MILAN trained on Places365 and CLIP-Dissect, using the spatial max summary function for CLIP-Dissect as well. Our results are generally mixed, with DnD performing the best (average of 0.7766), then MILAN (average of 0.7698), and finally CLIP-Dissect (average of 0.7390) using CLIP cosine similarity. These results match with BERTScore, as using that metric DnD performs the best (average of 0.8489), then MILAN (average of 0.8472), and finally CLIP-Dissect (average of 0.8368). With mpnet cosine similarity, CLIP-Dissect is calculated as the best (average of 0.1970) followed by MILAN (average of 0.1391) and then DnD (average of 0.1089). However, the MILANNOTATIONS dataset is very noisy as seen in Figure 14 and thus is largely unreliable for evaluation. The dataset provides three annotations per neuron, each made by a different person. This causes them to often be very different from one another, not representing the correct concept of the neuron. We can see in the table that the average CLIP cosine similarity between the annotations for each of these reliable neurons is 0.7884, the average mpnet cosine similarity is 0.1215, and the average BERTScore is 0.8482. Due to this noisiness, we found that the best performing description is as generic as possible as it will be closest to all descriptions. In fact, we found that a trivial baseline describing each neuron as ”depictions” irrespective of the neuron’s function outperforms all description methods, including MILAN. This leads us to believe MILANNOTATIONS should not be relied upon to evaluate description quality. Figure 14 showcases descriptions and MILANNOTATIONS for some neurons.
@@
Table 7: Similarity between predicted labels and ”ground truth” MILANNOTATIONS on the 4 intermediate layers of ResNet-152 trained on ImageNet. We observe that when using CLIP cosine similarity and BERTScore, DnD  performs the best followed by MILAN and then CLIP-Dissect, but when using MPNet cosine similarity, CLIP-Dissect performs the best followed by MILAN and then DnD. Simply labeling every neuron as ”depictions” outperforms all other methods, demonstrating the unreliability of MILANNOTATIONS as an evaluation method. The set’s noisiness is also shown by the average similarities between the different annotations for each neuron, demonstrating that the annotations are not consistent. We round results to the nearest 4 decimal places, but in instances of ties within a row, we observe further digits and bold accordingly.
@@
Metric Layer MILAN CLIP-Dissect Describe-and-Dissect (Ours) ”depiction” Avg Cos Sim Between Annotations

### v2

@@ Appendix A Appendix
In the appendix, we discuss specifics of the methods and experiments presented in the paper and provide additional results. Appendix Section A.1 provides further details on the method pipeline, Section A.2 presents additional qualitative evaluations, Section A.3 shows results for quantitative comparison to MILANNOTATIONS, Section A.4 provides detailed ablation studies on each step of the DnD pipeline, Section A.5 extends our land coverage use case experiment, Section A.6 presents an additional use for DnD as an OOD classifier, Section A.7 shows an extension of DnD to return multiple concept labels, Section A.8 provides an analysis of the diversity of our method’s labels, Section A.9 showcases failure cases of DnD, Section A.10 discusses the limitations and directions for future work, and Section A.11 outlines the broader impact of our work.
@@ A.1.3 GPT Prompt
      A.1.4 Scoring Function Details
@@
      A.1.5 Selecting a Scoring Function
    6. (f)
      A.1.6 Amazon Mechanical Turk Setup
@@ A.4.4 Multimodal GPT
      A.4.5 Effects of Large-Language-Model Choice
    6. (f)
      A.4.6 Effects of GPT Concept Summarization
@@ A.5 Use Case: Fixed Classifier
    A.6 Additional Use Case
  7. 7.
    A.7 Multiple Labels
  8. 8.
    A.8 Diversity Analysis
  9. 9.
    A.9 Failure Cases
  10. 10.
    A.10 Limitations
  11. 11.
    A.11 Broader Impact
@@ A.1.1 Detailed Description of Attention Cropping
The attention cropping algorithm described in Section 3.1 is composed of three primary procedures:
@@
  • Step 1. Global Thresholding using Otsu’s Method. For an activation map of neuron $n$, we define regions with high activations as the "foreground" and regions with low activations as the "background". Otsu’s Method (Otsu, 1979) automatically calculates threshold values which maximizes interclass variance between foreground and background pixels. Interclass variance $\sigma^{2}_{B}$ is defined by $\sigma^{2}_{B}=W_{b}W_{f}(\mu_{b}-\mu_{f})^{2}$, where $W_{b,f}$ denotes weights of background and foreground pixels and $\mu_{b,f}$ denotes the mean intensity of background and foreground pixels respectively. For global thresholding used in DnD, $W_{f}$ is the fraction of pixels in activation map $M_{n}(x_{i})$ above potential threshold $\lambda$ while $W_{b}$ is the fraction of pixels below $\lambda$. A value of $\lambda$ that maximizes $\sigma^{2}_{B}$ is used as the global threshold.
@@
  • Step 3. Crop Activating Images. Bounding boxes recorded from Step 2 are overlaid on corresponding activating images from $D_{probe}$, which are cropped to form $D_{cropped}$. To prevent overlapping attention crops, the IoU score between every crop in $D_{cropped}$ is less than an empirically set parameter $\eta$, where high $\eta$ values yield attention crops with greater similarities. We use $\eta=4$ across all experiments.
@@
Figure 5: Detailed Visualization of Attention Cropping Pipeline. All three steps of attention cropping are shown for Layer 2 Neuron 165. Steps 1 and 2 illustrate the derivation of bounding boxes from salient regions in the original activation map and are overlaid on original activating images from $\mathcal{D}_{probe}$ in Step 3. Cropped images are added back to $\mathcal{D}_{probe}$ to form $\mathcal{D}_{probe}\cup\mathcal{D}_{cropped}$.
A.1.2 Holistic vs. Local Concepts
To improve the accuracy of DnD, we form the probing dataset with cropped images representing local concepts and uncropped images representing holistic concepts. We define local concepts as concepts contained to a portion of the image and holistic concepts as emergent concepts represented by the whole image. This is necessary as solely using image crops can results in vague or spuriously correlated descriptions. We show two examples in Figure 6. In Layer 1 Neuron 64, uncropped images help produce nuanced concepts: cropped images only represent “blue” while uncropped images reflect the “aquatic” concept. Layer 3 Neuron 120 presents an example where cropped images activate on highly localized regions, making the image-to-text model prone to produce spurious captions. Uncropped images (last two images) make the “canine” concept more evident.
((a))
((b))
Figure 6: Holistic vs Local Concepts. We show two cases where a union of local and holistic concepts helps improve model accuracy. Augmenting the dataset reduces vague and spurious captioning.
A.1.3 GPT Prompt
The Concept Selection process utilizes GPT to identify and coherently summarize similarities between image captions of every image in $I$. Due to the generative nature of DnD, engineering a precise prompt is crucial to the method’s overall performance. The use of GPT in DnD can be split into two primary cases, each with an individual prompt: 1) Summarizing image captions to better bring out the underlying concepts, and 2) Detecting similarities between the image captions. We explain each part separately.
@@ A.1.3 GPT Prompt
"Only state your answer without a period and quotation marks. Do not number your answer. State one coherent and concise concept label that simplifies the following description and deletes any unnecessary details:"
@@
"Only state your answer without a period and quotation marks and do not simply repeat the descriptions. State one coherent and concise concept label that is 1-5 words long and can semantically summarize and represent most, not necessarily all, of the conceptual similarities in the following descriptions:"
In addition to the prompt, we use few shot-prompting to feed GPT two sets of example descriptions along with the respective human-identified similarities:
Example 1:
Human-Identified Similarity: "multicolored textiles"
Image Captions:
  • "a purple background with a very soft texture"
  • "a brown background with a diagonal pattern of lines and lines"
  • "a white windmill with a red door and a red door in the middle of the picture"
  • "a beige background with a rough texture of linen"
  • "a beige background with a rough texture and a very soft texture"
Example 2:
Human-Identified Similarity: "red-themed scenes"
Image Captions:
  • "a little girl is sitting in a red tractor with the word sofy on the front"
  • "a toy car sits on a red ottoman in a play room"
  • "a red dress with silver studs and a silver belt"
  • "a red chevrolet camaro is on display at a car show"
  • "a red spool of a cable with the word red on it"
We also include the prompt used for our Use Case one-shot-classification.
3. Use Case Superclass Classification. To classify neurons into broader NAIP superclasses we use the following prompt:
"State one coherent and concise concept label 1-5 words long related to landscapes/satellite imagery that semantically summarizes and represents most, not necessarily all, of the conceptual similarities in the following descriptions. Focus on colors, textures, and patterns. After, print one most likely natural landscape described by the satellite imagery captions, in parenthesis, on the same line. Be confident and do not be vague: "
A.1.4 Scoring Function Details
Though scoring functions all seek to accomplish the same purpose, the logic behind them can vary greatly and have different characteristics. Simplest scoring functions, such as mean, can be easily skewed by outliers in $\mathcal{H}_{j}$, resulting in final concepts that are loosely related to the features detected by neurons. In this section, we discuss three different scoring functions and a combination of them which we experimented with in Section A.1.5. The concept with highest score is chosen as the final description for neuron $n$.
@@ A.1.4 Scoring Function Details
    $\textit{score}_{M}(\mathcal{H}_{j})=-\frac{1}{Q}{\sum\limits_{i=1}^{Q}\mathcal{H}_{ji}}$
    $\textit{score}_{TK}(\mathcal{H}_{j},\beta)=-\frac{1}{\beta}\sum\limits_{i=1}^{\beta}\mathcal{H}_{ji}^{2}$
  • Image Products. Let set $\mathcal{D}_{j}^{t}\subset\mathcal{D}_{j}$, such that it keeps the $t$ highest activating images from $\mathcal{D}_{j}$. From the original set of activating images $I$ and $\mathcal{D}_{j}^{t}$, the Image Product score is defined as the average cosine similarity between original highly activating images and the generated images for the concept $j$. We measure this similarity using CLIP-ViT-B/16 Radford et al. (2021) as our image encoder $E(\cdot)$:
@@
    See Figure 7 for an illustration of Image Product. Intuitively, Image Products selects the candidate concept whose generated images are most similar to the original highly activating images. However, Image Product doesn’t actually account for how highly the new images activate the target neuron, which is why we chose to use this method to supplement other scoring functions. The enhanced scoring method, TopK Squared + Image Products, is described below.
  • TopK Squared + Image Products. This scoring function uses both image similarity and neuron activations to select the best concept. The method combines Image Products and TopK Squared by multiplying the Image Product score with the relative rankings of each concept’s TopK Squared score. We define $\mathcal{R}_{TK}=\{\ \textit{score}_{TK}(\mathcal{H}_{j},\beta),\;\forall\;j\in\{1,...,N\}\}$ as the set of TopK Squared scores for different descriptions. The final score is then:
    $\textit{score}_{TK-IP}(\mathcal{H}_{j},\beta,I,\mathcal{D}_{j}^{t})=\\ (N-\text{Rank}(\textit{score}_{TK}(\mathcal{H}_{j},\beta);\mathcal{R}_{TK}))\cdot\textit{score}_{IP}(I,\mathcal{D}_{j}^{t}),$
@@
A.1.5 Selecting a Scoring Function
We again use ResNet-50 with Imagenet $\cup$ Broden as the probing dataset. 50 neurons were randomly chosen from each of the 4 hidden layers, with each neuron evaluated twice, again rating the quality of descriptions on a scale 1-5. The participants in this experiment were volunteers with no knowledge of which descriptions were provided by which methods. Table 8 shows the performance of different scoring functions described in Section A.1.4. We observe that all three scoring functions presented perform similarly well, and shows our algorithm is robust and not dependent on the choice of scoring function.
@@ A.1.5 Selecting a Scoring Function
A.1.6 Amazon Mechanical Turk Setup
In this section we explain the specifics of our Amazon Mechanical Turk experiment setup in Section 4.3. In the interface shown in Figure 8, we display the top 10 highly activating images for a neuron of interest and prompt workers to answer the question: "Does description accurately describe most of the above 10 images?" Each given description is rated on a 1-5 scale with 1 representing "Strongly Disagree", 3 representing "Neither Agree nor Disagree", and 5 representing "Strongly Agree". Evaluators are finally asked to select the label that best describes the set of 10 images.
We used the final question to ensure the validity of participants’ assessments by ensuring their choice for best description is consistent with description ratings. If the label selected as best description in the final question was not (one of) the highest rated descriptions by the user, we deem the response invalid. In our analysis we only kept valid responses, which amounted to 1744/2400 total responses for Resnet-50 and 443/600 responses for ResNet-18 (around 75% of the responses were valid).
@@ A.1.6 Amazon Mechanical Turk Setup
One downside of our experimental setup is that it is not clear which highly activating images we should display to the users and how to display them, although these choices may be important to the final results. In our experiments, we displayed the images in $\mathcal{D}_{\text{probe}}$ with the highest mean activation as done by CLIP-Dissect (Oikarinen & Weng, 2023), but displaying images with highest max activation would likely be better for MILAN (Hernandez et al., 2022). Unfortunately there is no simple answer to which images we should display that would be good for all methods. See Limitations A.10 for more discussion.
@@ A.2 Qualitative Evaluation
Figures 9, 10, 11, and 12 provide supplementary qualitative results for DnD along with results produced by baseline methods for ResNet-50 and ResNet-18. Figure 13 showcases our descriptions on random neurons of the ViT-B/16 Vision Transformer trained on ImageNet, showing good descriptions overall and higher quality than those of CLIP-Dissect. Figure 14 extends DnD to ResNet-152, demonstrating the ability to scale to larger vision networks. In Figures 15 and 16 we dissect ResNet-50 using the CIFAR-100 training image set as $D_{probe}$, showcasing our method still perfoms well with a different probing dataset. Finally, we present a qualitative comparison with FALCON (Kalibhat et al., 2023) in Figure 17.
@@
Figure 14: Examples of DnD results from ResNet-152 (ImageNet). We showcase a set of randomly selected neurons and their descriptions across all 4 layers of ResNet-152 trained on ImageNet. DnD is scalable to larger vision networks.
@@
Figure 17: Examples of DnD results compared with FALCON results from layers across ResNet-50 (ImageNet). We showcase a set of randomly selected neurons provided in the appendix of Kalibhat et al. (2023) and their descriptions from layers across ResNet-50 trained on ImageNet-1K. DnD performs well compared to FALCON and other contemporary methods.
@@ A.3 Quantitative Results: MILANNOTATIONS
In this section we discuss using the MILANNOTATIONS dataset (Hernandez et al., 2022) as a set of ground truths to calculate quantitative results on the intermediate layers of ResNet-152. Table 9 displays an experiment done on 103 randomly chosen "reliable" neurons across the 4 intermediate layers of ResNet-152. A neuron was deemed "reliable" if, out of its three corresponding MILANNOTATIONS, two had a CLIP cosine similarity exceeding a certain threshold (for the purposes of this experiment we set that threshold to 0.81). Additionally, we define our summary function $g$ to be spatial max, as this is what was used to calculate the highest activating images for the MILANNOTATIONS. We compare against MILAN trained on Places365 and CLIP-Dissect, using the spatial max summary function for CLIP-Dissect as well. Our results are generally mixed, with DnD performing the best (average of 0.7766), then MILAN (average of 0.7698), and finally CLIP-Dissect (average of 0.7390) using CLIP cosine similarity. These results match with BERTScore, as using that metric DnD performs the best (average of 0.8489), then MILAN (average of 0.8472), and finally CLIP-Dissect (average of 0.8368). With mpnet cosine similarity, CLIP-Dissect is calculated as the best (average of 0.1970) followed by MILAN (average of 0.1391) and then DnD (average of 0.1089). However, the MILANNOTATIONS dataset is very noisy as seen in Figure 18 and thus is largely unreliable for evaluation. The dataset provides three annotations per neuron, each made by a different person. This causes them to often be very different from one another, not representing the correct concept of the neuron. We can see in the table that the average CLIP cosine similarity between the annotations for each of these reliable neurons is 0.7884, the average mpnet cosine similarity is 0.1215, and the average BERTScore is 0.8482. Due to this noisiness, we found that the best performing description is as generic as possible as it will be closest to all descriptions. In fact, we found that a trivial baseline describing each neuron as "depictions" irrespective of the neuron’s function outperforms all description methods, including MILAN. This leads us to believe MILANNOTATIONS should not be relied upon to evaluate description quality. Figure 18 showcases descriptions and MILANNOTATIONS for some neurons.
@@
Table 9: Similarity between predicted labels and "ground truth" MILANNOTATIONS on the 4 intermediate layers of ResNet-152 trained on ImageNet. We observe that when using CLIP cosine similarity and BERTScore, DnD  performs the best followed by MILAN and then CLIP-Dissect, but when using MPNet cosine similarity, CLIP-Dissect performs the best followed by MILAN and then DnD. Simply labeling every neuron as "depictions" outperforms all other methods, demonstrating the unreliability of MILANNOTATIONS as an evaluation method. The set’s noisiness is also shown by the average similarities between the different annotations for each neuron, demonstrating that the annotations are not consistent. We round results to the nearest 4 decimal places, but in instances of ties within a row, we observe further digits and bold accordingly.
@@
Metric Layer MILAN CLIP-Dissect Describe-and-Dissect (Ours) "depiction" Avg Cos Sim Between Annotations